# KnowCoder-X: Boosting Multilingual Information Extraction via Code

### v2

@@ Abstract
Empirical evidence indicates that LLMs exhibit spontaneous cross-lingual alignment. However, although LLMs show promising cross-lingual alignment in IE, a significant imbalance across languages persists, highlighting an underlying deficiency. To address this, we propose KnowCoder-X, a powerful code LLM with advanced cross-lingual and multilingual capabilities for universal information extraction. Firstly, it standardizes the representation of multilingual schemas using Python classes, ensuring a consistent ontology across different languages. Then, IE across languages is formulated as a unified code generation task. Secondly, we enhance the model’s cross-lingual transferability through IE cross-lingual alignment instruction tuning on a translated instance prediction task we proposed. During this phase, we also construct a high-quality and diverse bilingual IE parallel dataset with $\mathbf{257k}$ samples, called ParallelNER, synthesized by our proposed robust three-stage pipeline, with manual annotation to ensure quality. Although without training in $\mathbf{29}$ unseen languages, KnowCoder-X surpasses ChatGPT by $\mathbf{30.17\%}$ and SoTA by $\mathbf{20.03\%}$, thereby demonstrating superior cross-lingual IE capabilities. Comprehensive evaluations on $\mathbf{64}$ IE benchmarks in Chinese and English under various settings demonstrate that KnowCoder-X significantly enhances cross-lingual IE transfer through boosting the IE alignment. Our code and dataset are available at: https://github.com/ICT-GoKnow/KnowCoder.
@@
Yuxin Zuo1thanks:   Co-first authors., Wenxuan Jiang211footnotemark: 1, Wenxuan Liu [^1] footnotemark: 1, Zixuan Li1, Long Bai1, Hanbin Wang3, Yutao Zeng1, Xiaolong Jin1thanks:   Corresponding author., Jiafeng Guo1, Xueqi Cheng1 1Key Laboratory of Network Data Science and Technology, Institute of Computing Technology, Chinese Academy of Sciences 2School of Software, Northeastern University 3School of Software, Peking University {zuoyuxin24s, liuwenxuan2024z, jinxiaolong}@ict.ac.cn
@@ 1 Introduction
Cross-lingual Information Extraction (IE) focuses on automatically extracting structured information from texts in unseen languages, following manually designed schemas. Cross-lingual schema understanding and representation have long been a challenge. Large Language Models (LLMs), trained on massive multilingual corpora, have significantly advanced multilingual language processing (Brown et al., 2020; Touvron et al., 2023a, b). Previous studies (Zhang et al., 2023; Qi et al., 2023; Wang et al., 2024; Gao et al., 2024; Li et al., 2024a) have shown that LLMs exhibit spontaneous cross-lingual alignment to facilitate the transfer of abilities and knowledge across languages. Our findings suggest the presence of this alignment in IE, indicating a strong potential for improving the IE cross-lingual transfer. We first define IE parallel data, which is across various languages that share the same schema, sentences, and extracted instances. Previous studies used label projection to generate such data to alleviate the challenges of IE in low-resource Kolluru et al. (2022); Hennig et al. (2023); Rios et al. (2024). We construct an IE parallel dataset in Chinese based on CoNLL 2003 Sang and Meulder (2003) manually, to evaluate the LLaMA2-7B trained on English Named Entity Recognition (NER) datasets used in KnowCoder Li et al. (2024b). Table 1 shows the results. After training on English IE datasets, there was a notable enhancement in the parallel dataset of Chinese IE ($5.9$$\rightarrow$52.752.752.752.7), strongly supporting the presence of IE cross-lingual alignment. However, the significant performance gaps between the two languages ($95.1$ vs $52.7$) indicate that the alignment remains weak. To address this, we propose two strategies to enhance IE cross-lingual alignment in LLMs by aligning the schema and extraction process across different languages, respectively:
@@ 4 IE Cross-Lingual Alignment
KnowCoder-X adopts a two-phase instruction tuning framework. The first phase is the IE cross-lingual alignment phase, where we propose the task and data. Subsequently, we conduct instruction tuning on Chinese and English UIE tasks with $46$ IE datasets to obtain KnowCoder-X. Appendix H shows examples of instruction tuning data. Appendix G shows the detailed dataset statistics.
In this section, we introduce the IE cross-lingual alignment phase. First, we introduce the instruction and completion of the translated spans prediction task (§4.1). To obtain high-quality IE parallel data for this training phase, we propose a three-stage automatic pipeline for constructing IE parallel data (§4.2) and creating ParallelNER (§4.3).
@@ 6.1 Cross-Lingual Evaluation
To evaluate whether KnowCoder-X improves IE cross-lingual alignment, we conduct the cross-lingual evaluation on $9$ unseen languages of Multiconer22. We have also supplemented the evaluation in Chinese and English. Table 2 shows results. KnowCoder-X significantly outperforms ChatGPT and SoTA in the cross-lingual setting, achieving a $30.17\%$ and $20.03\%$ improvement, respectively. For $4$ unseen languages used in B2NER that constitute more than $0.1\%$ of the general LLM pretraining corpus Touvron et al. (2023c), KnowCoder-X outperforms B2NER by $19.36\%$, with results comparable to B2NER in both Chinese and English, demonstrating a significant improvement in cross-lingual transfer. Surprisingly, despite no training data in $9$ unseen languages, KnowCoder-X has achieved performance comparable to the supervised method in Spanish and Turkish that significantly outperforms the SoTA.
@@ 6.2 Supervised Evaluation
The results for NER, RE, EE (including ED and EAE) tasks are shown in Tables 3, 4, and 5 respectively. KnowCoder-X outperforms the multilingual SoTA baselines on most benchmarks for four tasks and ranks within the top-$2$ results across all RE and ED benchmarks. In the English IE, KnowCoder-X has achieved significant average improvements of $3.03$ and $2.92$ F1 points on the RE and ED tasks, with an improvement of $7.85$ points on the kbp37 of RE. In the Chinese IE, KnowCoder-X has consistently achieved SoTA, with $4.12$ points average improvement over the SoTA baseline on the EAE task. Moreover, it can be observed that the model not only achieves significant improvements in NER but also exhibits even greater enhancements in RE, ED, and EAE. This indicates that the alignment phase enhances the multilingual NER capability through cross-lingual alignment, which improves the performance of all tasks by leveraging the foundational spotting capabilities of NER.
@@ 6.4 Analysis on Cross-Lingual Transfer
7 Conclusion
We introduce KnowCoder-X, a code-based multilingual IE model that significantly enhances cross-lingual transfer through unified schema representation and IE cross-lingual alignment training. KnowCoder-X utilizes Python classes to represent schemas uniformly across languages, ensuring semantic consistency in multilingual IE. Additionally, the IE cross-lingual alignment phase leverages high-quality IE parallel datasets ParallelNER, which are generated by our proposed LLM-based pipeline, to boost multilingual generalization. Comprehensive experiments demonstrate that KnowCoder-X achieves SoTA in Chinese and English IE under various settings, with remarkable cross-lingual capability.
@@ 7 Conclusion
Expanding training to include languages beyond English and Chinese remains an area for future exploration. At the same time, incorporating additional languages would provide valuable insights into the cross-lingual alignment mechanisms within multilingual IE systems. Additionally, addressing domain specificity will be a key focus of future research. Specialized vocabulary and domain-specific language use present significant challenges for cross-lingual transfer, making it essential to overcome these obstacles in upcoming work. Introducing parallel data from other IE tasks during the cross-lingual alignment phase also offers promising avenues for further investigation.
@@ Appendix A Description Generation
  • Description Polish: Based on the description generated by the first phase, this phase further instructs the LLM to polish the description based on a set of entities iteratively. The details are as follows: 1) Randomly sample $20$ entities from the training set for iterative refinement. 2) Iterate over each entity from the entity set above: Format the current entity with the current description into the description polish template as shown in Table 21, instructing the LLM to evaluate whether the current description adequately covers the example semantically: If the current description adequately covers the entity, the LLM will not change the description. If the current description fails to encompass the entity, the LLM will revise the inappropriate parts of the description, ensuring the description covers the current entity.
@@ Appendix E Full Comparison of Supervised Evaluation
Appendix F Computational Resource Comparison
Appendix G Data Statistics
G.1 ParallelNER
G.2 Dataset Statistics

### v3

@@ Abstract
Empirical evidence indicates that LLMs exhibit spontaneous cross-lingual alignment. However, although LLMs show promising cross-lingual alignment in Information Extraction (IE), a significant imbalance across languages persists, highlighting an underlying deficiency. To address this, we propose KnowCoder-X, a powerful code LLM with advanced cross-lingual and multilingual capabilities for universal IE. Firstly, it standardizes the representation of multilingual schemas using Python classes, ensuring a consistent ontology across different languages. Then, IE across languages is formulated as a unified code generation task. Secondly, we conduct IE cross-lingual alignment instruction tuning on the translated instance prediction task to enhance the model’s cross-lingual transferability. During this phase, we also construct a high-quality and diverse bilingual IE parallel dataset with $\mathbf{257k}$ samples, called ParallelNER, synthesized by our proposed robust three-stage pipeline, with manual annotation to ensure quality. Although without training in $\mathbf{29}$ unseen languages, KnowCoder-X surpasses ChatGPT by $\mathbf{30.17\%}$ and SoTA by $\mathbf{20.03\%}$, thereby demonstrating superior cross-lingual IE capabilities. Comprehensive evaluations on $\mathbf{64}$ IE benchmarks in Chinese and English under various settings demonstrate that KnowCoder-X significantly enhances cross-lingual IE transfer through boosting the IE alignment. Our code and dataset are available at: https://github.com/ICT-GoKnow/KnowCoder.
@@
Yuxin Zuo1,2,3thanks:   Co-first authors., Wenxuan Jiang411footnotemark: 1, Wenxuan Liu1,2,311footnotemark: 1, Zixuan Li1,2thanks:   Corresponding authors., Long Bai1,2, Hanbin Wang5, Yutao Zeng1, Xiaolong Jin1,2,322footnotemark: 2, Jiafeng Guo1,2,3, Xueqi Cheng1,2,3 1Key Laboratory of Network Data Science and Technology, Institute of Computing Technology, Chinese Academy of Sciences 2State Key Laboratory of AI Safety 3School of Computer Science, University of Chinese Academy of Sciences 4School of Software, Northeastern University 5School of Software, Peking University {zuoyuxin24s, liuwenxuan2024z, lizixuan24h, jinxiaolong}@ict.ac.cn
@@ 1 Introduction
Cross-lingual Information Extraction (IE) focuses on automatically extracting structured information from texts in unseen languages, following manually designed schemas. Cross-lingual schema understanding and representation have long been a challenge. Large Language Models (LLMs), trained on massive multilingual corpora, have significantly advanced multilingual language processing (Brown et al., 2020; Touvron et al., 2023). Previous studies (Qi et al., 2023; Wang et al., 2024; Gao et al., 2024; Li et al., 2024a) have shown that LLMs exhibit spontaneous cross-lingual alignment to facilitate the transfer of abilities and knowledge across languages. Our findings suggest the presence of this alignment in IE, indicating a strong potential for improving the IE cross-lingual transfer. We first define IE parallel data, which is across various languages that share the same schema, sentences, and extracted instances. Previous studies used label projection to generate such data to alleviate the challenges of IE in low-resource Kolluru et al. (2022); Hennig et al. (2023); Rios et al. (2024). We construct an IE parallel dataset in Chinese based on CoNLL 2003 Sang and Meulder (2003) manually, to evaluate the LLaMA2-7B trained on English Named Entity Recognition (NER) datasets used in KnowCoder Li et al. (2024b). Table 1 shows the results. After training on English IE datasets, there was a notable enhancement in the parallel dataset of Chinese IE ($5.9$$\rightarrow$52.752.752.752.7), strongly supporting the presence of IE cross-lingual alignment. However, the significant performance gaps between the two languages ($95.1$ vs $52.7$) indicate that the alignment remains weak. To address this, we propose two strategies to enhance IE cross-lingual alignment in LLMs by aligning the schema and extraction process across different languages, respectively:
@@ 4 IE Cross-Lingual Alignment
KnowCoder-X adopts a two-phase instruction tuning framework. The first phase is the IE cross-lingual alignment phase, where we propose the task and data. Subsequently, we conduct instruction tuning on Chinese and English UIE tasks with $46$ IE datasets to obtain KnowCoder-X. Appendix H shows examples of instruction tuning data. Appendix G shows the detailed dataset statistics. In this section, we introduce the IE cross-lingual alignment phase. First, we introduce the instruction and completion of the translated spans prediction task (§4.1). To obtain high-quality IE parallel data for this training phase, we propose a three-stage automatic pipeline for constructing IE parallel data (§4.2) and creating ParallelNER (§4.3).
@@ 6.1 Cross-Lingual Evaluation
To evaluate whether KnowCoder-X improves IE cross-lingual alignment, we conduct the cross-lingual evaluation on $9$ unseen languages of Multiconer22. We have also supplemented the evaluation in Chinese and English. Table 2 shows results. KnowCoder-X significantly outperforms ChatGPT and SoTA in the cross-lingual setting, achieving a $30.17\%$ and $20.03\%$ improvement, respectively. For $4$ unseen languages used in B2NER that constitute more than $0.1\%$ of the general LLM pretraining corpus Touvron et al. (2023), KnowCoder-X outperforms B2NER by $19.36\%$, with results comparable to B2NER in both Chinese and English, demonstrating a significant improvement in cross-lingual transfer. Surprisingly, despite no training data in $9$ unseen languages, KnowCoder-X has achieved performance comparable to the supervised method in Spanish and Turkish that significantly outperforms the SoTA.
@@ 6.2 Supervised Evaluation
The results for NER, RE, EE (including ED and EAE) tasks are shown in Tables 3, 4, and 5 respectively. KnowCoder-X outperforms the SoTA baselines on most benchmarks for four tasks and ranks within the top-$2$ results across all RE and ED benchmarks. In the English IE, KnowCoder-X has achieved significant average improvements of $3.03$ and $2.92$ F1 points on the RE and ED tasks, with an improvement of $7.85$ points on the kbp37 of RE. In the Chinese IE, KnowCoder-X has consistently achieved SoTA, with $4.12$ points average improvement over the SoTA baseline on the EAE task. Moreover, it can be observed that the model not only achieves significant improvements in NER but also exhibits even greater enhancements in RE, ED, and EAE. This indicates that the alignment phase enhances the multilingual NER capability through cross-lingual alignment, which improves the performance of all tasks by leveraging the foundational spotting capabilities of NER.
@@ 6.4 Analysis on Cross-Lingual Transfer
7 Conclusions
We introduced KnowCoder-X, a code-based multilingual IE model that significantly enhances cross-lingual transfer through unified schema representation and IE cross-lingual alignment training. KnowCoder-X utilizes Python classes to represent schemas uniformly across languages, ensuring semantic consistency in multilingual IE. Additionally, the IE cross-lingual alignment phase leverages high-quality IE parallel datasets ParallelNER, which are generated by our proposed LLM-based pipeline, to boost multilingual generalization. Comprehensive experiments demonstrate that KnowCoder-X achieves SoTA in Chinese and English IE under various settings, with remarkable cross-lingual capability.
@@ 7 Conclusions
Expanding training to include languages beyond English and Chinese remains an area for further exploration. Moreover, incorporating additional languages would provide valuable insights into the cross-lingual alignment mechanisms. Additionally, addressing domain specificity will be a key focus of future research. Specialized vocabulary and domain-specific language use present significant challenges for cross-lingual transfer, making it essential to overcome these obstacles in upcoming work. Introducing parallel data from other IE tasks during the cross-lingual alignment phase also offers promising avenues for further investigation.
Acknowledgments
This work is partially funded by the National Natural Science Foundation of China under grants 62306299 and 62441229, the Lenovo-CAS Joint Lab Youth Scientist Project, the project under Grants No. JCKY2022130C039, and the Strategic Priority Research Program of the CAS under Grants No. XDB0680102. We thank anonymous reviewers for their insightful comments and suggestions.
@@ Appendix A Description Generation
  • Description Polish: Based on the description generated by the first phase, this phase further instructs the LLM to polish the description based on a set of entities iteratively.
@@ Appendix E Full Comparison of Supervised Evaluation
Appendix F Computational Resource Comparison
Appendix G Data Statistics
G.1 ParallelNER
G.2 Dataset Statistics